OpenStudy (anonymous):

Find the intervals in which the function is increasing and decreasing. Also, find and classify all the critical points of the function using the First Derivative Test. A(t)=t^5-5t^4+20t^3-17 I am having trouble setting the derivative equal to zero. Please help!

OpenStudy (amistre64):

increaseing or decreasing is a 2nd derivative test... not a 1st derivative test :)

OpenStudy (anonymous):

No...don't think so

OpenStudy (amistre64):

A' = 5t^4 -20t^3 +60t^2 A' = 5t^2(t^2 -4t +12) = 0

OpenStudy (amistre64):

12..6.2 (t-6) (t+2) t = 6, -2, and 0 are your crits

OpenStudy (amistre64):

what were you having issues with?

OpenStudy (amistre64):

slope is a function of the 1st derivative; concavity is a function of the 2nd derivative

OpenStudy (anonymous):

6 x -2 is -12, not +12

OpenStudy (amistre64):

hmmm.... you might be right :)

OpenStudy (amistre64):

my eyes were deceiving me... ;)

OpenStudy (amistre64):

if its got imaginary roots, then we probably dont need to both with it.... right? sqrt(16 - 48) = sqrt(-32) no "real" roots for this, go with t=0 as the critical, and it probably an inflection point

OpenStudy (anonymous):

Right. Thanks!